A stone is dropped from rest into a well. The sound of the splash is heard exactly 2.30 s later. Find the depth of the well if the air temperature is 6.0°C. m
Answer #1


speed of sound v = 331 + (0.6*T)


v = 331 + (0.6*6)


v = 334.6 m/s

for the stone


displacement y = -h

initial velocity voy = 0

acceleration ay = -g

y = voy*t + (1/2)*ay*t1^2

-h = -(1/2)*g*t1^2

h = (1/2)*g*t1^2


h= (1/2)*9.8*t1^2


h = 4.9*t1^2 .......( 1 )


for the sound


distance travelled h = v*t2

h = 334.6*t2 ............( 2 )


given


t1 + t2 = 2.3 s


from 1 and 2


4.9*t1^2 = 334.6*t2

4.9*t1^2 = 334.6*(2.3-t1)


t1 = 2.23 s


t2 = 2.3 - 2.23 = 0.07 s

depth d = 334.6*0.07 = 23.422 m